OpenStudy (anonymous):

find the sum of the given infinite geometric series. a =-8,r=-0.02 1 a.-7.84 b.-8.16 c.-7.8 d.-58

OpenStudy (anonymous):

Sum of an Infinite Geometric Sequence is given by: \[\large Sum = \frac{a}{r-1}\] Here a = -8 and r = -0.02.. Plug in the values and find the Sum..

OpenStudy (anonymous):

ok got it,it was 7.84 thank you !

OpenStudy (anonymous):

\[Sum = \frac{-8}{-0.02 - 1} = \frac{8}{1.02} = 7.84\] Yes you are right..